# CompdVision: Combining Near-Field 3D Visual and Tactile Sensing Using a Compact Compound-Eye Imaging System

### 12 Dec 2023 v1

@@ Abstract
As automation technologies advance, the need for compact and multi-modal sensors in robotic applications is growing. To address this demand, we introduce CompdVision, a novel sensor that combines near-field 3D visual and tactile sensing. This sensor, with dimensions of 22$\times$14$\times$14 mm, leverages the compound eye imaging system to achieve a compact form factor without compromising its dual modalities. CompdVision utilizes two types of vision units to meet diverse sensing requirements. Stereo units with far-focus lenses can see through the transparent elastomer, facilitating depth estimation beyond the contact surface, while tactile units with near-focus lenses track the movement of markers embedded in the elastomer to obtain contact deformation. Experimental results validate the sensor’s superior performance in 3D visual and tactile sensing. The sensor demonstrates effective depth estimation within a 70mm range from its surface. Additionally, it registers high accuracy in tangential and normal force measurements. The dual modalities and compact design make the sensor a versatile tool for complex robotic tasks. Videos are available at https://hkust-orisys.github.io/CompdVision.
@@ I Introduction
3D visual and tactile sensing are critical components in robotic manipulation. Traditional approaches often employ tactile sensors in conjunction with external cameras, which are either placed within the local environment [1][2] or mounted on the robotic arm’s wrist [3][4]. However, these setups have limitations. For instance, most depth cameras are relatively bulky and optimized for detecting objects at a distance, which is not ideal for close-up interactions. This distance-centric design often leads to occlusion problems, where the camera’s view is blocked, hindering accurate object localization when the gripper moves near the target.
@@
Figure 1: (a) Human thumb next to CompdVision. (b) CompdVision (red) is mounted on a Robotiq 2F-140 Gripper. (c) The compound-eye imaging system utilizes side stereo units with far-focus lenses for depth estimation, while the central tactile units, equipped with near-focus lenses, reduce external noise to enable precise tracking of marker movements for tactile sensing.
Vision-based tactile sensors such as GelSight [5] and Digit [6] have been increasingly utilized due to their capacity to capture detailed tactile information, proving beneficial in diverse applications [7][8]. These sensors typically use internal cameras to capture deformations in soft skin, thereby providing tactile data. Despite their advantages, many of these sensors are still quite large, limiting their applicability in compact systems. To address this problem, researchers have begun to explore the use of compound-eye imaging systems [9]. Inspired by the eyes of certain insects, these systems consist of vision units, each capturing a partial view of the visual field [10]. This design allows for a wide field of view and large area tactile perception while maintaining a compact form factor.
In this paper, we introduce CompdVision, a multi-modal sensor that leverages a compact compound-eye imaging system, combining 3D visual and tactile sensing into a single unit. The sensor’s micro-scale vision units facilitate near-field 3D visual sensing within a 70mm range from the sensor surface, ensuring precise object localization throughout the grasping process. Post-contact, the sensor also provides a detailed contact deformation field, offering a comprehensive sensing solution for complex robotic tasks.
@@
The CompdVision sensor is equipped with a 3$\times$5 grid of vision units, as shown in Fig. 1. This design strategy not only expands the field of view but also offers the flexibility to assign each vision unit to different modalities. The vision units at the top-left and bottom-left corners constitute the left stereo pair, while the units at the top-right and bottom-right corners form the right stereo pair. Two pairs are dedicated to stereo depth estimation. Equipped with far-focus lenses, they are specialized in seeing through the transparent elastomer and capturing visual data, thereby facilitating robust depth estimation beyond the contact surface. Simultaneously, the remaining nine vision units, centrally located on the grid, are specifically designed to track the displacement of markers on the elastomer for tactile sensing. High-resolution images are derived by stitching together the partial views captured by these tactile units. These units are equipped with near-focus lenses, which minimize noise interference from external scenes, thus enhancing their tactile sensing capabilities. The placement of stereo vision units ensures that depth estimation effectively covers the central area where tactile sensing occurs. This configuration allows for a seamless combination of 3D visual and tactile sensing, enhancing the sensor’s performance in complex robotic tasks.
@@
  • A compound-eye imaging system combines 3D visual and tactile sensing into a single sensor unit.
  • Achievement of near-field 3D visual sensing, facilitated by micro-scale vision units, effective within a 70mm range from the sensor surface.
  • A compact sensor that is easily integrated into robotic systems.
@@ II-A Vision-Based Tactile Sensor
In 2009, Johnson et al. [11] employed photometric stereo techniques to capture tactile information. Utilizing a single camera and a tri-color lighting setup, this sensor was able to reconstruct surface deformations with high accuracy and resolution. Building on this, Yuan et al. [5] further refined the sensor by incorporating dot markers to estimate contact forces precisely. The TacTip family of sensors [12] leveraged internal pins to track and record surface deformations, offering another innovative approach to tactile sensing. Lambeta et al. [6] showcased the Digit sensors, which were mounted on a multi-fingered robotic hand and demonstrated capabilities in manipulating glass marbles.
However, a common limitation across these pioneering efforts has been the relatively large form factor of the sensors. To tackle this challenge, the GelSlim sensor [13] was redesigned with an optical path that utilized light guides and mirror reflections to optimize finger thickness and tactile sensing area. To further reduce size while retaining functionality, some researchers have begun to incorporate multiple cameras into their designs. For instance, the OmniTact sensor [14] achieved finger-sized dimensions and captured multi-directional, high-resolution tactile information using an array of micro-cameras. Trueeb et al. [15] introduced another sensor, utilizing a multi-camera setup, featuring a large contact surface and a slim profile. Zhang et al. [10] presented a sensor that employed a novel compound eye imaging system comprised of an array of pinholes on a single complementary metal oxide semiconductor (CMOS) sensor, reducing the thickness to just 5 mm. Chen et al. [16] optimized the sensor by utilizing a microlens array-based vision system, maintaining the same slim profile while capturing clearer images.
@@ II-B Sensor Combining Visual and Tactile Sensing
Recent advancements in robotic manipulation have underscored the importance of combining vision-based tactile sensors with external cameras, inspiring researchers to integrate visual and tactile sensing capabilities into multi-modal sensors. Yamaguchi et al. [17] introduced the FingerVision sensor, featuring a transparent skin marked with dot patterns, and an internal RGB camera. The sensor captures tactile data by tracking dot marker displacements and visual data through the transparent elastomer. Hogan et al. [18] designed a semi-transparent skin sensor capable of functioning as a tactile sensor for high-resolution surface deformation or as a visual sensor, depending on the internal lighting conditions. This sensor was further developed to estimate object proximity [19]. The SpecTac sensor [20] employs UV LEDs and UV fluorescent markers on a transparent elastomer, enabling it to switch between modes for capturing high-quality visual and tactile data. Similarly, the TIRgel sensor [21] adjusts focus to alternate between visual and tactile modalities. Despite these innovations, these sensors are constrained by their use of single cameras, limiting their ability to accurately measure the distance to external objects. To address this, Shimonomura et al. [22] developed a sensor using an infrared camera for tactile sensing and dual cameras for stereo depth estimation. Despite these advancements, the sensor cannot provide contact details due to its rigid surface. Roberge et al. [23] further introduced the StereoTac sensor, incorporating a semi-transparent skin that can switch between tactile and visual sensing modes based on internal lighting conditions. However, the sensor faces limitations regarding image clarity due to the semi-transparent skin and the inability to operate both sensing modes simultaneously. Furthermore, the size of these multi-modal sensors remains challenging for compact integration, necessitating future research to address these concerns. In our work, we aim to address these limitations and enhance the capabilities of multi-modal sensors, particularly focusing on the simultaneous operation of sensing modalities and compact integration.
@@ III Sensor Design and Fabrication
The touching subsystem serves a dual purpose: providing the sensor’s structural foundation and enabling its tactile functionality. This subsystem features transparent elastomer layers and embedded dot markers. The transparency of these elastomer layers enables the stereo units of the imaging subsystem to capture external visual data. Meanwhile, the opaque black dot markers allow the tactile units to perceive the movement of these dots, which correlates with the deformation of the sensor surface. The touching subsystem houses the imaging subsystem and protects the entire assembly. The imaging subsystem interprets the visual and tactile data. Its compound-eye design accommodates the distinct requirements of the stereo and tactile units. The imaging subsystem’s design and functionality are crucial to the comprehensive performance of the sensor, contributing significantly to its ability to accurately detect and interpret sensing signals.
The touching subsystem’s fabrication begins by adhering an acrylic plate to the 3D-printed sensor shell using UV glue. Subsequently, Solaris (Smooth-On, Inc.), a transparent soft silicone, is poured into the structure, forming the base elastomer layer. Following the curing of Solaris, the sensing pattern is created by stamping silicon-based ink onto the base elastomer layer using a pillar array. The ink is composed of Psycho Paint, Smooth-On Silc-Pigment 9 black color paste, and silicon thinner mixed in a weight ratio of 4:12:3. A second layer of Solaris elastomer is then poured onto the mold and covered by the previous component, resulting in a protective elastomer layer that shields the dot markers.
@@
Figure 3: (a) depicts images captured by a stereo pair focused on a calibration chessboard. The visualization of the stereo calibration results, showing the positions of the chessboard and vision units, is illustrated in (b) for the left stereo pair and in (c) for the right stereo pair.
TABLE I: Stereo Calibration Results
@@
[…]
Figure 5: Scheme of tactile sensing: (a) An initial image is used to crop ROIs based on the centers of common markers. These cropped ROIs are then stitched together in the non-overlapping areas to form a high-resolution image. (b) Utilizing the same cropping regions, a stitched image is formed from post-contact captures. (c) The RGB image is converted to HSV color space and thresholded to isolate the markers, followed by using SimpleBlobDetector for marker localization. (d) The detected blob points are compared and matched with those from the initial image, establishing correspondences based on nearest neighbors. This step facilitates the mapping of marker displacements onto a regular grid, resulting in a 2D displacement field.
[…]
IV Sensing Principle
This section elaborates on the sensing principles of stereo and tactile units. These units form the core of our data acquisition system, with the stereo units providing 3D vision perception and the tactile units offering sensor surface interactions.
IV-A Stereo Units
Stereo vision is a technique that mimics human vision by extracting 3D information from a scene captured from two different viewpoints. This method leverages the disparity between corresponding points in the two images to calculate the depth of each point in the scene [24].
CompdVision sensor initiates the process by cropping the regions of interest from the raw image according to specified stereo units. Contrary to traditional designs, these units’ vertical alignment necessitates a preprocessing step to transpose the images, thereby producing left and right images compatible with conventional stereo image processing algorithms.
Before undertaking depth estimation, calibration is a crucial preparatory step to ensure precision. This phase is facilitated using a chessboard pattern, depicted in Fig. 3(a), in conjunction with the Stereo Camera Calibrator app in MATLAB [25]. This assists in acquiring crucial parameters such as the intrinsic parameters (including the focal length and optical centers), extrinsic parameters (representing the geometric relationship between the cameras), lens distortion coefficients, and a transformation matrix vital for aligning the two images. The visualization of the calibration results for both stereo units is illustrated in Fig. 3(b) and (c). The specific values for the rectified focal length, baseline, and the reprojection error are detailed in Table I.
The depth estimation process, as illustrated in Fig. 4, commences by extracting the image tiles from the specified stereo units. The tiles are then transposed to align horizontally, followed by distortion correction to rectify any irregularities. The tiles are then converted to grayscale to prepare for the application of the Semi-Global Block Matching (SGBM) algorithm [26][27]. The SGBM algorithm is chosen for its balance of accuracy and computational efficiency. Subsequently, a right disparity map is generated, which is crucial for the application of Weighted Least Squares (WLS) filtering. WLS filtering refines the disparity map, reducing noise and improving image quality. Finally, the refined disparity map is obtained. Any values below the minimum disparity are set to NaN, which helps handle ambiguous regions, such as repetitive or texture-less areas, where block-matching algorithms struggle to identify unique matches. This measure ensures more accurate and reliable depth estimation. Moreover, any values in the marker area of the disparity map are removed, as the depth of markers does not contribute to our study.
IV-B Tactile Units
@@ IV-B1 Image Stitching
The process begins with identifying Regions of Interest (ROIs) within the initial image, selected for its distinct white background that simplifies blob detection and facilitates accurate determination of the centers of common markers. These centers provide a reference for defining the ROIs. The ROIs are chosen to exclude overlaps between units and to remove black and unutilized regions, as shown in Fig. 5(a), ensuring that each unit’s data is distinct and effectively utilized. After the identification of the ROIs, they are cropped from the initial image. Because of the inversion of the image projected onto the CMOS sensor, each cropped tile is flipped along both the X and Y axes. These adjusted ROIs are then stitched together, producing a high-resolution pre-contact image. When force is applied, a similar process is adopted using the same cropping regions to produce a stitched post-contact image, as depicted in Fig. 5(b). This approach guarantees consistency and accuracy, thereby enabling reliable comparisons between the pre-contact and post-contact states.

### 15 Mar 2024 v2

@@ Abstract
As automation technologies advance, the need for compact and multi-modal sensors in robotic applications is growing. To address this demand, we introduce CompdVision, a novel sensor that employs a compound-eye imaging system to combine near-field 3D visual and tactile sensing within a compact form factor. CompdVision utilizes two types of vision units to address diverse sensing needs, eliminating the need for complex modality conversion. Stereo units with far-focus lenses can see through the transparent elastomer for depth estimation beyond the contact surface. Simultaneously, tactile units with near-focus lenses track the movement of markers embedded in the elastomer to obtain contact deformation. Experimental results validate the sensor’s superior performance in 3D visual and tactile sensing, proving its capability for reliable external object depth estimation and precise measurement of tangential and normal contact forces. The dual modalities and compact design make the sensor a versatile tool for robotic manipulation.
@@ I Introduction
3D visual and tactile sensing significantly advance robotic manipulation by using depth cameras and tactile sensors, enhancing the robot’s capabilities in dynamic environments [1][2]. This fusion provides depth perception, object recognition and contact information, essential for executing robotic tasks with high autonomy and adaptability [3][4].
@@
Figure 1: (a) Human thumb next to CompdVision. (b) CompdVision (red) is mounted on a gripper. (c) The compound-eye imaging system utilizes side stereo units with far-focus lenses for depth estimation, while the central tactile units, equipped with near-focus lenses, reduce external noise to enable precise tracking of marker movements for tactile sensing.
Vision-based tactile sensors, revolutionizing robotic touch perception by translating physical interactions into visual data [5], have proven invaluable across diverse applications [6][7]. Miniaturization efforts, notably through the compound-eye imaging system [8], have achieved sensor sizes akin to human fingertips [9][10]. Although focused on tactile sensing, the system’s design, where each vision unit operates independently and is customizable, paves the way for developing compact and multi-modal sensors.
Using cameras and tactile sensors separately has been a common approach in robotics. A single sensor combining visual and tactile sensing offers a significant advancement, improving robots’ interaction and perception capabilities [11]. However, this innovation encounters some challenges. The reliance on 2D visual sensing limits depth perception and spatial understanding. Low-rate conversion between visual and tactile modalities [12][13], presents challenges in dynamic manipulation tasks [14]. Furthermore, achieving a compact sensor combining 3D visual and tactile sensing is difficult due to bulky camera modules and the considerable spacing between stereo cameras. Such spacing results in non-overlapping fields of view when objects are close to cameras, necessitating the placement of the sensor surface centimeters away from cameras to perceive depth for close-up interactions [15]. Addressing these challenges is vital for enabling precise and adaptable robotic manipulation.
TABLE I: Comparison of CompdVision and other Sensors
Sensor Modality Modality conversion Sensing rate Size (in mm)
Vision
Tactile
GelSight[5]
No
Force + Shape
30fps
35$\times$35$\times$60
OmniTact[9]
No
Shape
30fps
$\varnothing$30$\times$33
Zhang et al. [10]
No
Force
30fps
14.5$\times$14.5$\times$5
FingerVision[11]
RGB, limited
Force
Simultaneous sensing
30fps
47$\times$40$\times$30
Finger-STS[14]
RGB
Marker Flow
LEDs
NA
50$\times$50
SpecTac[12]
RGB
Force
LEDs
0.2s visual-tactile cycle
NA
StereoTac[15]
RGBD
Shape
LEDs
4Hz LEDs alternating rate
NA
CompdVision (Ours)
RGBD
Force
Simultaneous sensing
25fps
22$\times$14$\times$14
In this paper, we introduce CompdVision, a sensor that leverages a compound-eye imaging system to combine near-field 3D visual and tactile sensing within a compact form factor. Before making contact, it collects visual and depth information, and after contact, it offers deformation field and force measurements. The sensor’s compact size, along with its ability to perform 3D visual and tactile sensing simultaneously without the need for modality conversion, advances the field of robotic sensing.
@@
The CompdVision sensor is equipped with a 3$\times$5 grid of vision units, as shown in Fig. 1. The vision units at the top-left and bottom-left corners constitute the left stereo pair, while the units at the top-right and bottom-right corners form the right stereo pair. Equipped with far-focus lenses, two pairs are dedicated to seeing through the transparent elastomer and capturing visual data, thereby facilitating robust depth estimation beyond the contact surface. Simultaneously, the remaining nine vision units, centrally located on the grid, are designed to track the displacement of markers on the elastomer for tactile sensing. The partial views captured by each tactile unit are stitched, expanding the field of view. These units are equipped with near-focus lenses, which minimize noise interference from external scenes for enhanced tactile sensing capabilities. The placement of stereo vision units ensures that depth estimation effectively covers the central area where tactile sensing occurs.
@@
  • Employ a compound-eye imaging system for a compact and dual-modal sensor.
  • Achievement of 3D visual sensing with micro-scale vision units for near-field perception.
  • Simultaneous visual and tactile sensing using a dual-focus microlens array, eliminating modality conversion.
@@ II-A Vision-Based Tactile Sensor
In 2009, Johnson et al. [16] introduced a tactile sensor using photometric stereo techniques and a tri-color lighting setup to reconstruct surface deformations. Yuan et al. [5] enhanced this by adding dot markers for force estimation. The TacTip sensors [17] used internal pins for deformation tracking. Lambeta et al. [18] developed Digit sensors for a robotic hand, enabling intricate object manipulation.
Despite this progress, the sensor size remained a challenge. The GelSlim sensor [19] employed light guides and mirrors for a slimmer profile. The OmniTact sensor [9] and Trueeb et al.’s design [20] achieved compact dimensions with a multi-camera setup, pushing the boundaries of sensor miniaturization. However, these attempts faced challenges due to the inherent bulkiness of camera modules, even when employing fish-eye or micro-cameras. Zhang et al. [10] innovated by integrating micro-scale vision units into a compound-eye imaging system, significantly slimming sensor thickness to 5 mm. Chen et al. [21] further optimized this with a microlens array, enabling high-resolution imaging without internal lighting. While these sensors with compound eyes were focused on tactile sensing, their innovative design laid the groundwork for developing compact and multi-modal sensors. Each vision unit within the system operates independently and can be customized for specific needs, enabling the combination of diverse modalities.
@@ II-B Sensor Combining Visual and Tactile Sensing
FingerVision [11] demonstrates visual-tactile sensing with its transparent skin and dot markers. However, its internal RGB camera’s focus on markers for enhanced tracking restricts it to close object visualization. TIRgel [13] shows the conversion between tactile and visual modality via focus adjustment. Hogan et al. [22] developed a sensor with semi-transparent skin that operates as either a tactile or visual sensor depending on internal lighting conditions. Such modality conversion limits it to switching to tactile sensing only after making contact [14]. SpecTac [12] utilizes UV LEDs and fluorescent markers on a transparent elastomer for visual-tactile sensing. The sensor alternates between modalities by flickering the UV LEDs in a duty cycle. This method of modality conversion leads to a low visual-tactile sensing rate.
Single-camera setups in these sensors limit depth estimation capabilities, crucial for precise external object interactions. Shimonomura et al. [23] introduced a sensor combining infrared and stereo cameras for tactile and depth sensing, respectively, though its rigid surface limits contact detail acquisition. StereoTac [15], featuring a semi-transparent skin, alternates between modes using LED lighting but faces challenges with low-rate modality conversion and struggles with compact integration due to the bulkiness of camera modules. Additionally, it shows a great challenge in designing compact sensors combining 3D visual and tactile sensing due to the millimeter-scale spacing between stereo cameras. This spacing prevents complete overlap of the cameras’ fields of view, particularly resulting in blind spots and impaired depth perception when objects are close. Consequently, the sensor surface must be positioned centimeters away from the cameras for close-up interactions, restricting the development of compact sensors. Furthermore, while some sensors can reconstruct the high-resolution shape of the sensor surface, they lack the capability for marker tracking or force estimation, essential for detecting slippage [24].
In comparison to existing sensors (partially summarized in Table I), we develop a compact, dual-modal sensor using a compound-eye imaging system. We achieve near-field 3D visual sensing with micro-scale vision units and simultaneous visual and tactile sensing through a dual-focus microlens array without the need for modality conversion.
@@ III Sensor Design and Fabrication
The touching subsystem, utilizing transparent elastomer with embedded dot markers, serves both structural and functional purposes, allowing for visual and tactile data collection. The transparency of the elastomer enables stereo units within the imaging subsystem to capture external visual data, while the tactile units track the markers’ movements to obtain surface deformations. The touching subsystem protects the entire assembly and houses the imaging subsystem, which is crucial in processing both visual and tactile inputs. The compound-eye design of the imaging subsystem is customized to fulfill the different requirements on focal length for stereo and tactile vision units, ensuring the sensor’s ability to capture and interpret 3D visual and tactile information.
@@
The touching subsystem’s fabrication begins by adhering an acrylic plate to the 3D-printed sensor shell using UV glue. Subsequently, Solaris (Smooth-On, Inc.), a transparent soft silicone, is poured into the structure, forming the base elastomer layer. Following the curing of Solaris, the sensing pattern is created by stamping silicone-based ink onto the base elastomer layer using a 3D-printed pillar array. The ink is composed of Psycho Paint, Smooth-On Silc-Pigment 9 black color paste, and silicone thinner mixed in a weight ratio of 8:6:3. A second layer of Solaris elastomer is then poured onto the mold and covered by the previous component, resulting in a protective elastomer layer that shields the dot markers. The total thickness of the elastomer is about 6mm.
@@
[…]
Figure 4: Scheme of tactile sensing: (a) An initial image is used to crop ROIs based on the centers of common markers. These cropped ROIs are then stitched together in the non-overlapping areas to form a single image. (b) Utilizing the same cropping regions, a stitched image is formed after contact. (c) The RGB image is converted to HSV color space and thresholded to isolate the markers, followed by using SimpleBlobDetector for marker localization. (d) The detected blob points are compared and matched with those from the initial image, establishing correspondences based on nearest neighbors.
IV Sensing Principle
This section details the sensing principles of the 3D visual and tactile modalities. The 3D visual modality enables depth perception through the sensor’s surface, while the tactile modality captures surface deformation. We also present the simultaneous sensing of these two modalities.
IV-A 3D Visual Modality
Stereo vision, a technique mimicking human vision, extracts 3D information from scenes captured from differing viewpoints [25]. By positioning vision units at opposite corners, our sensor forms stereo pairs, ensuring sufficient overlap in their fields of view for effective stereo matching, especially when objects come into contact with the sensor. This configuration also guarantees high-resolution depth perception over extended distances. Furthermore, the placement of two pairs on the sides ensures coverage of the central area, where tactile sensing occurs.
Stereo calibration, conducted with a chessboard pattern using MATLAB’s Stereo Camera Calibrator [26], emphasizes our sensor’s small baseline (the distance between vision units), crucial for enabling near-field 3D visual sensing capabilities. Unlike traditional stereo vision approaches that might struggle with near-field depth estimation due to larger baselines, our sensor’s compact design significantly reduces blind spots in close proximity. Consequently, the sensor surface can be positioned as close as about 11mm from the imaging system, marking an advancement in compact sensor design. Calibration results are detailed in Table II.
TABLE II: Stereo Calibration Results
@@ IV Sensing Principle
As illustrated in Fig. 3, the depth estimation process begins by extracting image tiles from the specified stereo pair. The vertical arrangement of these units, unlike traditional setups, necessitates the transposition of tiles to align left and right images for standard stereo processing. This is followed by distortion correction to rectify any irregularities. The tiles are then converted to grayscale in preparation for applying the Semi-Global Block Matching (SGBM) algorithm [27][28]. Subsequently, a right disparity map is generated, which is essential for applying Weighted Least Squares (WLS) filtering. WLS filtering enhances the disparity map by reducing noise and improving image quality. Finally, the refined disparity map is obtained. Any values below the minimum disparity are set to NaN, which helps handle ambiguous regions, such as repetitive or texture-less areas, where matching algorithms struggle to identify unique matches. Furthermore, any values within the marker area of the disparity map are excluded, as the depth of markers does not contribute to our study.
IV-B Tactile Modality
@@ IV-B1 Image Stitching
As shown in Fig. 4(a), the central vision unit captures eight markers, including two centrally located and six near the edges. Upon contact, there’s a risk that edge markers move out of view and become untrackable by a single unit. To address this, surrounding units can be used to expand the field of view, ensuring all markers are tracked with sufficient margins for post-contact movement.
The process starts with capturing an initial image, featuring a clear white background without any contact on the sensor surface. This background simplifies blob detection and aids in determining the centers of common markers, which serve as references for defining Regions of Interest (ROIs). Selected ROIs prevent overlaps between units and exclude black and unused areas, ensuring each unit’s view is unique and effectively utilized. After identifying the ROIs, they are cropped and stitched into a single image. Upon applying force, the same cropping regions produce a stitched post-contact image, as shown in Fig. 4(b). This stitched image ensures the tracking of all markers after contact.